Can you please help me out with this? This is very difficult for me. I understand that the pattern between the numbers is that you are squaring the number each time, but that only makes it harder for me to solve. I don't know where to begin. Can you please explain how solve this problem? I would REALLY appreciate it!!
What is the sum of the geometric sequence 4, 16, 64, … if there are 8 terms?

Answer by richwmiller(17219)

It has two patterns
2^2, 4^2, 8^2 which is not geometric
and
4, 4*4, 4*16 which is geometric,
Your ratio (r) is 4
t1=4 n=8 r=4
Tn = t1 * r^(n - 1)
S=t1*(1 - r^n)/(1 - r)
S=4*(1 - 4^8)/(1 - 2)
S = 262140